Dear Richard dear all,

please find below a short summary of the activities done today on the mountain related to the AO system. Italian people working at LBTO were:

Arcetri: E. Pinna, A. Puglisi, L. Fini, M. Sozzi, P. Stefanini, S. Esposito, C. Arcidiacono
ADS: E. Anaclerio

W unit

1) Functionality test of the unit has been repeated after the new of the short power outage happened last night. Result was positive and no new problem has been found.
2) The WFS board has been installed on the W stages using the reference pin to try to keep the alignment obtained in Arcetri.
3) The cabling of the board has been connected to all devices.
4) A problem was found in the BCU39 behavior. The unit is working on and off. A similar problem was encountered in Arcetri when the connector of the 3.3 power supply of the BCU was not properly in place.
5) A complete functionality test of the W unit has been done after re cabling. Two devices namely the CCD39 and the cameral lens translation stage has not been checked yet due to the BCU39 problem that does not allow to see the frames from the CCD39. All the other devices were working correctly.
6) A test of the stage motion under different rotation angle of the AGW (corresponding to different gravity loads) has been done after installation of the WFS board. The test was successful.

AO control SW

1) Fixed bug in AdSec Arbitrator code. Submitted to SVN@Arcetri.
2) Writing a procedure to compare directory trees on local and/or remote machines. It'll be used as a tool to maintain alignment among various installations.
3) Installation procedure revised to separate installation of executables from calibration and configuration data
4) A lap top computer was configured with the AO control SW to be used as control computer during the planned main computers stop of Thursday.

LBT672a

1) The clean tent in the optical lab has been slightly moved and plastic adjusted in size to have complete air circulation.
2) The cooling lines for the secondary has been positioned in place.
3) The LBT672a unit has been transferred in the optical lab. The unit has been unpacked and placed under the clean tent.
4) The electronic cabinet has been transferred in the optical lab.

Next update tomorrow.

All the best

Simone

-- ArmandoRiccardi - 02 Apr 2010
Topic revision: r1 - 02 Apr 2010, ArmandoRiccardi
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback